### v3

# Interstellar comet 3I/ATLAS: discovery and physical description

Abstract
We present the characteristics of interstellar comet 3I/ATLAS, discovered on 2025 July 1 by the Asteroid Terrestrial-impact Last Alert System. The comet has eccentricity 6.08 and velocity at infinity 57 km/s, indicating an interstellar origin. We obtained B,V, R, I, g, r, i, and z photometry at Kottamia Astronomical Observatory, Palomar Observatory, and Apache Point Observatory on 2025 July 2, 3, and 6. We measured colour indices B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05, r–i=0.160.03, i–z=–0.020.07, g–i=1.000.05, and a spectral slope of 16.01.9 /100 nm. We estimate the comet’s dust cross-section within 10,000 km to be 184.64.6 km2, assuming a 0.1 albedo. 3I/ATLAS’s coma has FWHM2.2 arcsec and A(0∘)f=280.83.2 cm. We estimate 3I/ATLAS is ejecting µm to mm dust at 0.01-1 m/s, with a mass loss of 0.1 - 1.0 kg/s.
keywords:
minor planets, comets: general1 Introduction
Interstellar objects (ISO) are small bodies that originate from outside the solar system whose orbital trajectories intersect the solar system (for a review, see Jewitt, 2024). Present-day studies of extra-solar debris disks are limited to their micron and mm-sized components (Su et al., 2024; Xie et al., 2025), but ISOs provide opportunities for the study of macroscopic bodies that originated from other stellar systems (e.g., Bolin et al., 2018; Hopkins et al., 2023). Until recently, only two known interstellar objects, 1I/‘Oumuamua and 2I/Borisov have been found (Jewitt, 2024). 1I/’Oumuamua appeared inactive but was subsequently found to have significant non-gravitational acceleration (Micheli et al., 2018), while 2I/Borisov showed clear evidence of an extended coma (Guzik et al., 2019; Kim et al., 2020; Bolin & Lisse, 2020) and volatile compounds (Bodewits et al., 2020; Cordiner et al., 2020).
The third known macroscopic interstellar object, comet 3I/ATLAS, was discovered on 2025 July 1 (Denneau et al., 2025) by the Asteroid Terrestrial-impact Last Alert System (ATLAS). This paper describes its discovery, and its subsequent imaging characterisation by the Kottamia Astronomical Observatory (KAO) 1.88-m telescope, the Palomar 200-inch telescope, and the Apache Point Observatory (APO) Astrophysical Research Consortium 3.5-m (ARC 3.5-,m) telescope. We use the techniques described by Bolin et al. (2020, 2024) for characterizing active objects with multi-band photometry to test the coma-centric colours of 3I/ATLAS. In addition, we apply the surface brightness profile analysis techniques of Bolin et al. (2021, 2025a) to characterize the activity of 3I/ATLAS and its dust properties.
2 Observations
Initially designated as A11pl3Z, interstellar comet 3I/ATLAS was discovered at the ATLAS Rio Hurtado, Chile facility (Minor Planet Center (MPC) observing code W68). The initial discovery observations of 3I/ATLAS were made on 2025 July 1 05:15:11 in the ATLAS o-band filter as seen in Panels a-d of Figure 1 (Denneau et al., 2025). The comet was discovered at right ascension, declination of 18:07:27.68 and -18:41:40.2 and was detected in four o-band exposures taken between 2025 July 1 05:15:11and 2025 July 1 06:20:31. A list containing details of the ATLAS observations is presented in Section S.1 and Table S1 of the Supplemental Material.
For the follow up and characterization of 3I/ATLAS, used the Kottamia Astronomical Observatory 1.88-m telescope’s (KAO 1.88-m’s) Kottamia Faint Imaging Spectro-Polarimeter (KFISP), the Next Generation Palomar Spectrograph (NGPS) instrument on the Palomar 200-inch telescope, and the Apache Point Observatory Astrophysical Research Consortium 3.5-m (ARC 3.5-m)/ Astrophysical Research Consortium Telescope Imaging Camera (ARCTIC) on 2025 July 2, 2025 July 3, and 2025 July 6. Imaging in B,V,R, and I, was taken with KFISP, g, r, and i images were taken with the NGPS, and g, r, i, and z images were taken with ARCTIC. Spectra of 3I/ATLAS were also taken with P200 and ARC 3.5-m, which will be described in a forthcoming paper (Belyakov et al., 2025b).
Images from all three facilities were reduced using biases and flats taken for each of the filters. Image stacks of the NGPS and ARCTIC data are shown in the mosaic images in Figs. 2 and 3. The multi-band data were calibrated using the Pan-STARRS1 catalog (Chambers et al., 2016) and the colour transformations from (Tonry et al., 2012) to convert Pan-STARR1 photometry magnitudes into SDSS photometry magnitudes. The photometry of 3I/ATLAS was measured using a 4″ radius circular aperture (10,000 km from the comet’s nucleus at its 3.36 - 3.45 au distance from the Earth during our KAO 1.88-m, P200, and ARCTIC 3.5-m observaitons). We used a 6.5-7.8″ radius sky background annulus to measure our photomery. The seeing at the KAO 1.88-m, the P200, and the ARC 3.5-m measured between 1.3-1.6 arcsec in images containing the comet. A list of the details of the KAO 1.88-m, P200, and ARC 3.5-m observations is presented in Section S.1 and Table S1 of the Supplemental Material.
3 Results
3.1 Orbit
The astrometry of 3I/ATLAS from ATLAS taken in the discovery images on 2025 July 1 was reported to the MPC. Astrometry of the comet in 319 observations taken from 97 observational facilities around the world between 2025 June 14.25 UTC and 2025 July 4 UTC was subsequently reported to the MPC 111https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=3I, accessed on 2025 July 4.. The orbital solution from JPL HORIZONS, using observations taken between June 14, 2025, and July 4, 2025, is shown in Table S2. The location and orbit of 3I/ATLAS at the time of its discovery on 2025 July 1 is shown on Figure 4 based on values from JPL Horizons222https://ssd.jpl.nasa.gov/horizons/app.html#/.
3.2 Visible photometry
We computed the brightness of 3I/ATLAS in the KAO 1.88-m images of B=18.90 0.21, V=17.92 0.01, R=17.22 0.05, and I=17.08 0.09, in the NGPS images of g=18.720.02, r=17.860.01, and i=17.720.02, and g=18.580.06, r=17.770.02, i=17.590.04, and z=17.610.05 in the ARCTIC images. We used the NGPS r-band stack to search for evidence of an extended coma in 3I/ATLAS. The surface brightness profile extracted from the r-band images of 3I/ATLAS (blue line in Panel D of Figure 2) shows evidence of being extended, having 2-10 times the flux compared to the surface brightness profile of a comparison star (orange line in the same panel) outside of 1.3 arcsec of both profiles. The FWHM of the coma in the NGPS r-band stack is 2.2 arcsec compared to the 1.6 arcsec FWHM of background stars in the sidereal-tracked images. 3I/ATLAS ’s colours are B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05 mag, r–i=0.160.03 mag, i–z=–0.020.07 mag, and g–i=1.000.05 mag corresponding to a spectral slope of 16.01.9 /100 nm., somewhat more red than interstellar comet 2I/Borisov which had g–r=0.630.05, r–i=0.200.02, g–i=0.830.05, and a spectral slope of 9.71.8 /100 nm. The colors of 3I/ATLAS derived from our observations are similar to those reported by other photometric and spectroscopic studies (e.g., Opitom et al., 2025). We plot the g–r, r–i, and r–z colours of 3I/ATLAS in Figure 5.
We use the g, r, and i photometry obtained by the P200/NGPS to calculate the A(0∘)f parameter, a proxy for dust production rate normalised to a phase angle of 0 degrees (A’Hearn et al., 1984; Schleicher et al., 1998; Marcus, 2007). We find g-band A(0∘)f = 187.74.1 cm, r-band A(0∘)f = 280.83.1 cm, i-band A(0∘)f = 287.84.5 cm, and z-band A(0∘)f = 289.313.7 cm, typical values for active solar system comets (A’Hearn et al., 1995; Kelley et al., 2013). We calculate the absolute magnitude, , according to the following equation (Jewitt, 1991):
where is the apparent V magnitude (determined from our g and r observations to be = 18.280.03), is the heliocentric distance of the comet on 2025 July 3 of 4.43 au, is the distance of the comet from the Earth on 2025 July 3 of 3.43 au, is a function describing the brightening of the comet which we approximate with = 0.04, appropriate for comets observed at phase angles with 20 degrees (Bertini et al., 2017), and is the phase angle of the comet in degrees, of 2.65 degrees. Using Eq. 1, we obtain =12.270.03 mag, cautioning that the uncertainties of inferred from Eq. 1 are systematically underestimated due to the presently unknown phase function for 3I/ATLAS. We translate the magnitude computed from Eq. 1 into an effective cross-section, , in units of km2 within a 10,000 km aperture using
from Jewitt et al. (2016), where is the albedo of the comet, assumed to be 0.10, typical for comet dust (Jewitt & Meech, 1986; Kolokolova et al., 2004; Zubko et al., 2017), finding a value of 184.64.6 km2.
We can use the size of the coma measured in the sunward direction to provide an estimate of the dust ejection velocity, , (Jewitt & Meech, 1987):
where Dt is the extent of the coma in the sunward direction measured in km, is a quantity inversely proportional to the particle size, , in µm, = 0.006 m s-2, and is the heliocentric distance in au. We calculate Dt using (Hsieh et al., 2021):
where is the half-width half maximum measured from our r-band stack of 1.1 arcsec, is the measured half-width half maximum of background stars of 0.8 arcsec, resulting in a of 0.7 arcsec, which translates to 1850 km at the 3.43 au distance of the comet. Plugging our calculated value of of 1850 km into Eq. 4 gives 1.1m s-1 suggesting that m-scale to mm-scale particles are ejected from the comet at 0.01-1 m/s. The dust mass loss, can be roughly estimated from , A(0∘)f, the dust albedo , the radius of the dust particles, , and the dust density, , with the following relation derived from (Fink & Rubin, 2012):
plugging in our values of = 10-2-1 m/s for µm to mm-sized particles, A(0∘)f = 281 cm, = 750 kg m-3 (Fulle et al., 2016), and = 0.1 results in = 0.1 - 1.0 kg/s, similar to interstellar comet 2I/Borisov when it was on its inbound trajectory and located 2.7 au from the Sun (Fitzsimmons et al., 2019; Jewitt & Luu, 2019), and similar to some Jupiter Family Comets in the solar system (Gillan et al., 2024).
4 Discussion and conclusion
Interstellar comet 3I/ATLAS exhibits an active appearance and colours similar to some solar system comets, as seen in Figures 2 and 5. However, it seems to be significantly more red in its color compared to the other interstellar comet 2I (g–r=0.630.05, r–i=0.200.05, g–i=0.830.05, and r–z=–0.030.06), however, is less red compared to 1I/’Oumuamua (g–r=0.840.05, r–i=0.310.09, g–i=1.150.10, and r–z=0.410.10). However, this latter comparison may be inaccurate due to the fact that 3I/ATLAS is an active body, and 1I/’Oumuamua was only very weakly active as evidenced by the visible lack of a coma (Micheli et al., 2018). A complete table of the physical properties of the different interstellar objects is included in Table S3. Interstellar comet 3I/ATLAS appears to be more red than the majority of comets as seen in Fig. 4. The increased redness of 3I compared to 2I and other active comets may be due to the nature of its activity if it lacks the emission of gas molecules present in the g-band wavelength such as CN, C2, and C2 (Farnham et al., 2000). With our small sample size of three objects, it appears that the interstellar objects span a wide range in color. Additionally, 3I shows evidence of having a similar dust mass loss compared to 2I, but has a much higher dust mass loss rate than 1I, which was observed to show a negligible amount of dust mass loss.
While it is difficult to speculate on the size of 3I due to the coma obscuring its nucleus, future high-resolution observations will provide a better test of its size than ground-based, seeing-limited observations can. Additional future observations of 3I/ATLAS in the visible will test its activity and potentially its evolution as it passes through perihelion in 2025 October at a distance 1.35 au from the Sun. Observations covering the near-infrared and submillimeter wavelengths will enable the study of its ice and dust grain properties. Space-based observations spanning the near-infrared to mid-infrared will facilitate a test of the production of cometary gases in its coma and activity-driving mechanism. Dedicated and serendipitous observations by surveys such as the Rubin Observatory Legacy Survey of Space and Time (Ivezić et al., 2019; Bolin et al., 2025b; Dorsey et al., 2025) will provide coverage of distant comets and ISOs which may lead to a greater understanding of the extrasolar planetesimal population (Hopkins et al., 2023).
Acknowledgements
The authors wish to acknowledge Larry Denneau for sharing the discovery images of 3I/ATLAS and the details of the discovery observations. The authors also wish to thank Robert Jedicke for help with the orbital diagram, and for the scientific discussion of the paper, and Yanga. R. Fernández for providing comments to the manuscript. Our work is based on observations taken at Palomar Observatory. The authors wish to thank the Palomar observing staff for their help with the observations. The authors also wish to recognize and acknowledge the cultural significance that Palomar Mountain has for the Pauma Band of the Luiseño Indians. MK and GJ acknowledge support from the NSF under grant AST-2205736 and the NASA under grants 80NSSC22K0479, 80NSSC24K0380, and 80NSSC24K0436. MWC acknowledges support from the National Science Foundation with grant numbers PHY-3117998, PHY-2308862, and PHY-2409481. MJG acknowledges support from the National Science Foundation with grant numbers PHY-2117997. The National Research Institute of Astronomy and Geophysics team acknowledges financial support from the Egyptian Science, Technology Innovation Funding Authority (STDF) under grant number 48102. Based on observations obtained with the Apache Point Observatory 3.5-meter telescope, which is owned and operated by the Astrophysical Research Consortium. MM thanks the projects MEYS LM2023032, LM2023047.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author. The data for the Palomar and Apache Point observations can be found at the following repository (Belyakov et al., 2025a).
Supplemental Material
The Supplemental Material for this manuscript is available online.
Supplemental Material
S.1 Observational details
Asteroid Terrestrial-impact Last Alert System Chile, Rio Hurtado: The ATLAS telescope consists of a 0.5-m reflector telescope equipped with a charge-coupled device (CCD) camera, which has a pixel scale of 1.86 arcseconds per pixel. Survey observations use an “orange” o-band custom filter providing wavelength coverage between 560 nm and 820 nm, and an effective wavelength of 663 nm (Tonry
et al., 2018). W68 is located at El Sauce Observatory, Rio Hurtado.
Kottamia Astronomical Observatory (KAO), Egypt, 1.88-m telescope/Kottamia Faint Imaging Spectro-Polarimeter (KFISP):
The KFISP is a combination direct imaging and spectrograph instrument mounted on the cassegrain focus of the KOA 1.88-m telescope (Osman, 2001; Azzam
et al., 2022). The KFISP in imaging mode consists of a 2048 pixel x 2048 pixel array with a 8 arcmin x 8 arcmin field of view and a 0.243 arcsec/pixel pixel scale. The imager possesses Johnson-Cousins B,V, R, and I filters (Cousins, 1976). In total, two 240 s B images, three 240 s V images, three 180 s R, and two 180 s I images were taken. The comet was tracked at a non-sidereal rate consistent with its skyplane motion over an airmass range of 1.50-1.56. The seeing measured in R-band images determined from full-width at half-maximum (FWHM) measurements of background stars of 1.33 arcsec during the 2025 July 2 observations. Kottamia Astronomical Observatory is located in the Cairo Governorate Desert, Egypt.
Palomar 200-inch telescope (P200)/Next Generation Palomar Spectrograph (NGPS) The NGPS instrument is mounted at cassegrain focus on the P200 and possesses a guide camera with a 4 arcmin x 4 arcmin field of view, 0.26 arcsec/pixel spatial scale (Tang
et al., 2023; Kasliwal
et al., 2024), and SDSS-equivalent g, r, and i filters ( = 467.2 nm/614.1 nm/745.8 nm, FWHM = 126.3 nm/115.0 nm/68.3 nm; Gunn
et al., 1998). The seeing in r-band was 1.6 arcsec measured by determining the full-width at half-maximum of background stars during the observations taken of 3I/ATLAS on 2025 July 3. Two g-band exposures and two i-band exposures were taken with a total cumulative exposure time of 150 s, and three r-band exposures were taken with a total cumulative exposure time of 270 s. The photometry from the g images was measured in individual images, and combined. Additionally, a sidereal-tracked image containing 3I/ATLAS was taken to provide star detections to compare with the surface brightness profile of 3I/ATLAS. The g, r, and i images were taken over an airmass range of 1.79 to 2.05 during the 2025 July 3 observations. The P200 is located on Palomar Mountain, California.
Apache Point Observatory Astrophysical Research Consortium 3.5-m (ARC 3.5-m)/ Astrophysical Research Consortium Telescope Imaging Camera (ARCTIC): ARCTIC has a 7.85 arcmin x 7.85 arcmin field of view and a 0.228 arcsec/pixel scale when binned in 2 x 2 mode (Huehnerhoff
et al., 2016). We conducted photometric time series observations of 3I/ATLAS using the ARCTIC imaging camera with the Sloan Digital Sky Survey (SDSS) g, r, i, and z filters with effective wavelengths 477 nm, 623 nm, 763 nm and 913 nm respectively (Fukugita et al., 1996). In total, 5 x 120 s g band images, 6 x 120 s r band images, 5 x 120 s i band images, and 4 x 120 z band images were taken. The ARC 3.5-m was tracked at the rate of motion of the comet, the seeing was measured to be 1.47 arcsec, and the airmass range of the observations was 2.30 to 1.81 during the 2025 July 6 observations. The ARC 3.5-m is located at Apache Point, New Mexico.
Facility | Instrument | UT Date and Time | r | Seeing4 | Airmass | Sky motion | m5 | ||
(au) | (au) | (∘) | (″) | (arcsec/h) | (mag) | ||||
W68 | ATLAS | 2025-07-01 05:15 | 3.50 | 4.50 | 2.0 | 2.0 | 1.05 | 73.3 | 17.83 |
KAO 1.88-m | KFISP | 2025-07-02 21:04 | 3.45 | 4.45 | 2.5 | 1.3 | 1.53 | 74.1 | 17.21 |
P200 | NGPS | 2025-07-03 07:27 | 3.43 | 4.43 | 2.6 | 1.6 | 1.63 | 75.8 | 17.86 |
ARC 3.5-m | ARCTIC | 2025-07-06 03:23 | 3.36 | 4.34 | 3.6 | 1.5 | 2.05 | 78.6 | 17.77 |
Notes. (1) Geocentric distance, (2) heliocentric distance, (3) phase angle, (4) measured in science images, (5) o/R/r magnitude as measured in the images taken during the observations.
Heliocentric Elements | |
Epoch (JD) | 2,460,859.5 |
Time of perihelion, (JD) | 2,460,636.9175586(3.76x10-5) |
Semi-major axis, (au) | –0.26549(2.49x10-3) |
Eccentricity, | 6.0823(9.76x10-2) |
Perihelion, (au) | 1.3493(1.33x10-2) |
Inclination, (∘) | 175.10961(4.87x10-3) |
Ascending node, (∘) | 322.08(0.11) |
Argument of perihelion, (∘) | 128.07(0.14) |
Mean Anomaly, (∘) | –854.68(10.65) |
name | q | e | i | g-r | r-i | r-z | g-i | Af | Refs. | |
(au) | (deg) | (cm) | (kg/s) | |||||||
1I/’Oumuamua | 0.26 | 1.20 | 122.74 | 0.840.05 | 0.360.1 | 0.360.10 | 1.200.11 | 0 | 0 | Meech et al. (2017) |
2I/Borisov | 2.01 | 3.36 | 44.05 | 0.630.05 | 0.200.02 | –0.030.11 | 0.830.05 | 14310 | 1 | Bolin et al. (2020) |
3I/ATLAS | 1.35 | 6.08 | 175.11 | 0.840.05 | 0.160.03 | 0.140.08 | 1.000.05 | 280.83.2 | 1 | This work. |
Notes. (1) perihelion distance, (2) eccentricity, (3) inclination, (4) g–r colour, (5) r–i colour, (6) g–i colour, (7) Af, (8) dust mass loss.